# A semi-parametric marginalized dynamic conditional correlation framework

### v3

@@ 5.2 Evaluation of VaR and ES forecasts
DCC-AL 67% 83% 81% 8% 42% 0
Lastly, Bayer and Dimitriadis (2022) propose three versions of ES backtests named as Auxiliary, Strict and Intercept ES regression (ESR) backtests [^5] For the implementation, we use the R package developed by the authors, which can be found at: https://search.r-project.org/CRAN/refmans/esback/html/esr_backtest.html. The link also includes the details of the three versions of the backtests. . These ESR backtests (two-sided), labeled as A, S, and I respectively, are also employed to backtest the ES forecasts from the 10 competing models.
As in Table 6, the three backtests return quite consistent results, with the DCC-AL model being not rejected. The models that get rejected at the 5% level by all three backtests are DCC-N, DCC-clik-N and DCC-t. Meanwhile, the semi-parametric DCC-QML and DCC-clik-QML do not get rejected.
@@
DCC-AL 87% 87% 71% 0
5.3 Minimum ES portfolios
Unlike univariate semi-parametric risk forecasting models, such as ES-CAViaR or CARE, the semi-parametric DCC-AL can be used for asset allocation. Therefore, in this section we evaluate the out-of-sample performance of the DCC-AL model in computing global minimum ES portfolios at the 2.5% level. We compare the properties of the minimum ES portfolios obtained by DCC-AL with those obtained by standard DCC models, which are used here as benchmarks. Our interest for the computation of minimum ES portfolios is motivated by the prominent role that this risk measure is assuming in the international regulatory framework.
In order to compute minimum ES portfolios, we develop a two-stage procedure. At the initial stage, an estimate of the conditional variance and covariance matrix $H_{t}$ is obtained by fitting a DCC-AL model to an equally weighted portfolios. The optimal portfolio weights are then identified by minimizing the portfolio ES conditional on the first stage volatility estimates. There are two main reasons for our preference for this type of approach: first, it is computationally convenient as it avoids re-estimating the model at each iteration, and second, it can be easily implemented for semi-parametric models other than the DCC-AL model, thus facilitating comparison with benchmark DCC models.
@@ 5.3 Minimum ES portfolios
    Given an initial set of portfolio weights $\mathbf{w}$, fit a DCC-AL (or other standard DCC models) and save $\hat{H}_{t}$ for each in-sample time stamp $t$
@@
For all models out-of-sample minimum ES portfolios are computed. Figure 5 illustrates the DCC-AL optimized minimum ES portfolio weights for all the forecasting steps. Table 7 compares the statistical properties of the minimum ES portfolio obtained by the DCC-AL model with the benchmarks. Firstly, we focus on the value of the empirical ES of optimized portfolios which is expected to be higher (closer to 0) for more effective risk minimization strategies. Compared to the EW portfolio, all estimated DCC models yield substantial reductions in terms of out-of-sample empirical ES. Meanwhile, DCC-AL and the other DCC models under analysis produce quite similar performances for the whole-out-of-sample period, with DCC-AL and DCC-clik-QML slightly outperforming the other DCC models. Such results lend evidence on the general validity of the proposed DCC-AL for asset allocation purposes. Regarding other descriptive statistics of portfolio returns, the DCC-AL produces smallest kurtosis and is least skewed. It also has the most conservative minimum value. These results provide evidence that the returns of the minimum ES portfolio obtained by the DCC-AL model are characterized by lighter tails and a less elongated left tail compared to the benchmarks, leading to a lower probability of occurrence of negative extreme events.
Figure 5: Boxplot of the DCC-AL minim ES portfolio weights for all forecasting steps.
Table 7: Minimum ES portfolio returns under different portfolio configurations produced by various models. Portfolio descriptive statistics include: $\widehat{ES}_{\alpha}$: sample $\alpha$-level ES ; $\kappa$: sample kurtosis; $skew$: sample skewness; $min$: sample minimum.
DCC-QML DCC-clik-QML DCC-t EW DCC-AL
@@
In addition, to examine the differences in the performance of the different models over a shorter time horizon, Figure 6 shows the moving empirical tail-mean (ES) of the optimised portfolio returns computed over a 250-day annual window. It is of particular interest to focus on the “COVID-19” period, which coincides with the negative peak in the plot, approximately centred on day 1000. First, we see that, in this highly risky period, the EW allocation is not effective in risk diversification since it yields remarkably lower ES values. At the same time, the DCC-t model, which generates competitive portfolio ES for the entire forecast period, is outperformed by all other semi-parametric DCC-type models for the high volatility “COVID-19” period, demonstrating the effectiveness of the semi-parametric approach, which is free of return distribution selection. Finally, focusing on the semi-parametric methods, it is evident that DCC-AL is able to produce preferred portfolio ES forecasts compared to DCC-QML and DCC-clik-QML for the “COVID-19” period.
@@
Figure 6: Moving ES (sample tail mean) of optimized portfolio returns.
6 Conclusions
In this paper, we propose an innovative framework for forecasting portfolio tail risk in a multivariate and semi-parametric setting. The proposed framework is capable of modeling high dimensional return series parsimoniously and efficiently. Compared to the state-of-the-art univariate semi-parametric models, our approach delivers competitive risk forecasting performances and, in addition, it can be used for a wider range of applications including asset allocation. Compared to existing DCC estimation approaches, our empirical application shows that the DCC-AL is able to deliver more accurate risk forecasts. Shifting the focus to the computation of minimum ES portfolios, we find that the DCC-AL produces competitive performances compared to other DCC models, especially for the highly volatile “COVID 19” period.
Potential projects for future research include a deeper investigation of the performance of the DCC-AL model in asset allocation, as well as an extension of the proposed framework to other settings such as the prediction of CoVaR (Adrian and Brunnermeier, 2016), building on recent results in semi-parametric estimation by Fissler and Hoga (2024) and Dimitriadis and Hoga (2024).
Finally, as our analysis in this paper is based on low-frequency daily data, our research plans also include extending the DCC-AL model by incorporating high-frequency realized measures to further improve its performance.
Disclosure of Interest
No conflict of interests to be declared.
Disclosure of Funding
Data Availability Statement
Appendix A
Simulation study
A simulation study has been conducted in order to assess the statistical properties of the two-step AL loss based estimation procedure discussed in Section 3. In this Appendix we provide a detailed description of the simulation design and of the main findings of our study. Our analysis has two main objectives. First, the study evaluates the bias and efficiency of the estimators of the DCC-AL parameters ($a,b,q,\gamma_{0}$). Second, we also assess the the risk forecasting performances of the estimated DCC-AL model in an equally weighted portfolio, via comparing the one-step-ahead $2.5\%$ level portfolio VaR and ES forecast accuracy to the “true” simulated values. In line with the empirical findings arising from our empirical application, we have considered the Data Generating Process (DGP) as a parametric DCC model of dimension $n$=28 with the true values of the correlation dynamic parameters in Equation (13) given by $a=0.12$ and $b=0.78$:
where $\Sigma_{\bm{\epsilon}}$ is a $n\times n$ matrix with unit diagonal and off-diagonal elements equal to 0.5. In the return equation $\mathbf{r}_{t}=\bm{\mu}_{t}+H^{1/2}_{t}\mathbf{z}_{t}$, the conditional mean vector $\bm{\mu}_{t}$ of returns is chosen to be zero. The DGP univariate volatilities are assumed to follow the GARCH(1,1) model:
$h^{2}_{t,i}=0.1+0.1r^{2}_{t-1,i}+0.8h^{2}_{t-1,i},\qquad i=1,\ldots,n.$
While the first two distributions belong to the spherical family, the same does not hold for the third one ($nst$). The DCC-AL model is then fitted to the time series of equally weighted $(\mathbf{w}=1/n)$ portfolio returns generated from the chosen DGP, for three different sample sizes $T\in\{2000,3000,5000\}$. Overall, matching 3 different distributional assumptions and sample sizes, we have 9 simulation settings and, under each of these, 250 return series have been generated. Conditional on the chosen error distribution and on the values of the correlation parameters $a$ and $b$, the true values of the other two parameters $q$ and $\gamma_{0}$ in the DCC-AL model can be then calculated as follows. In the spherical case, multivariate Normal and $t$, letting $F(.)$ be the CDF of $z_{t,i}$, $q=F^{-1}(\alpha)$ and $c=E(z_{t,i}|z_{t,i}\leq q)$. By Equation (15) it is then easy to obtain:
$\gamma_{0}=\log\left(\frac{c^{2}}{q^{2}}-1\right).$
Therefore, taking the multivariate Normal case as an example, in the DCC-AL the true value of $q^{2}$ is $\Phi^{-1}(\alpha)^{2}=(-1.96)^{2}=3.8415$. The true value for $c^{2}$ is $\left(\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}\right)^{2}=(-2.3378)^{2}$, where $\Phi$ is the standard Normal CDF and $\phi$ is standard Normal Probability Density Function (PDF). Further, we have $(1+\text{exp}(\gamma_{0}))=\frac{c^{2}}{q^{2}}=\frac{(-2.3378)^{2}}{(-1.9600)^{2}}=1.4227$, thus $\gamma_{0}=\text{log}(1.4227-1)=-0.8610$. These true values of $a$, $b$, $\gamma_{0}$ and $q^{2}$ for the multivariate Normal distribution are shown in the True rows in Table A1. When $\mathbf{z}_{t}\overset{iid}{\sim}t_{n}(\mathbf{0},I_{n};\nu=10)$, the true values of $q$, $c$ and $\gamma_{0}$ are obtained through a similar approach, replacing the analytical formulas for $q$ and $c$ with the equivalent expressions for an univariate $t$ distribution. In the $nst$ case, the true values of $q$, $c$ and $\gamma_{0}$ need to be calculated via simulations since this distribution is not a member of the spherical family. It also follows that, differently from what observed for the multivariate Normal and $t$ cases, the values of these coefficients will be dependent on the chosen portfolio allocation. So they could change when moving from equal weighting to a different allocation scheme. More specifically, $q$ and $c$ can be calculated by the following procedure.
  1. 1.
  2. 2.
    $z^{*}_{j}=\frac{r^{*}_{j}}{h^{*}_{j}},$
    for $j=1,\ldots,T_{sim}$ and where
  3. 3.
[…]
Table A1: DCC-AL model parameter estimates and VaR and ES forecasting results with simulated datasets.
DGP $a$ $b$ $\gamma_{0}$ $q^{2}$ $\widehat{Q}_{T+1}$ $\widehat{\text{ES}}_{T+1}$
$\mathcal{N}$, $T=2000$ True 0.1200 0.7800 -0.8610 3.8415 -1.3704 -1.6346
Mean 0.1360 0.7150 -0.9131 3.9308 -1.3849 -1.6420
RMSE 0.0751 0.1995 0.1639 0.2057 0.0606 0.0762
$\mathcal{N}$, $T=3000$ True 0.1200 0.7800 -0.8610 3.8415 -1.3565 -1.6180
Mean 0.1310 0.7319 -0.8919 3.9079 -1.3669 -1.6244
RMSE 0.0615 0.1736 0.1200 0.1613 0.0521 0.0611
$\mathcal{N}$, $T=5000$ True 0.1200 0.7800 -0.8610 3.8415 -1.3665 -1.6299
Mean 0.1265 0.7528 -0.8830 3.9018 -1.3760 -1.6367
RMSE 0.0466 0.1118 0.1029 0.1363 0.0392 0.0454
$t$, $T=2000$ True 0.1200 0.7800 -0.5097 3.9717 -1.4353 -1.8159
Mean 0.1591 0.6719 -0.5844 4.1136 -1.4479 -1.8108
RMSE 0.1023 0.2419 0.1894 0.2967 0.0811 0.1034
$t$, $T=3000$ True 0.1200 0.7800 -0.5097 3.9717 -1.3825 -1.7491
Mean 0.1497 0.6894 -0.5658 4.0686 -1.3990 -1.7543
RMSE 0.0884 0.2248 0.1527 0.2167 0.0692 0.0850
$t$, $T=5000$ True 0.1200 0.7800 -0.5097 3.9717 -1.4155 -1.7909
Mean 0.1303 0.7395 -0.5459 4.0619 -1.4304 -1.7995
RMSE 0.0617 0.1564 0.1172 0.1613 0.0534 0.0703
$nst$, $T=2000$ True 0.1200 0.7800 -0.8311 3.8772 -1.3570 -1.6259
Mean 0.1440 0.7030 -0.8893 3.9012 -1.3538 -1.6096
RMSE 0.0894 0.2245 0.1656 0.1849 0.0639 0.0762
$nst$, $T=3000$ True 0.1200 0.7800 -0.8311 3.8772 -1.3806 -1.6542
Mean 0.1356 0.7223 -0.8657 3.8978 -1.3799 -1.6461
RMSE 0.0729 0.1926 0.1246 0.1539 0.0576 0.0706
$nst$, $T=5000$ True 0.1200 0.7800 -0.8311 3.8772 -1.4001 -1.6775
Mean 0.1277 0.7486 -0.8569 3.9091 -1.4057 -1.6788
RMSE 0.0496 0.1195 0.1029 0.1251 0.0383 0.0465

### v4

@@ 5.2 Evaluation of VaR and ES forecasts
Semi-DCC-AL 67% 83% 81% 8% 42% 0
Semi-DCC-FZ0 81% 91% 80% 11% 40% 0
$\bm{\alpha=5\%}$
CARE-SAV 0% 1% 3% 1% 2% 5
CARE-IG 0% 1% 3% 1% 2% 5
ES-CAViaR-SAV 0% 0% 1% 0% 1% 5
ES-CAViaR-IG 0% 0% 2% 1% 3% 5
DCC-QML 0% 0% 0% 0% 0% 5
DCC-N 1% 1% 1% 0% 0% 5
DCC-clik-QML 0% 0% 1% 0% 0% 5
DCC-clik-N 1% 2% 4% 0% 0% 5
DCC-t 1% 1% 1% 0% 0% 5
Semi-DCC-AL 1% 2% 7% 0% 12% 3
Semi-DCC-FZ0 1% 2% 7% 0% 5% 3
Lastly, Bayer and Dimitriadis (2022) propose three versions of ES backtests named as Auxiliary, Strict and Intercept ES regression (ESR) backtests [^5] For the implementation, we use the R package developed by the authors, which can be found at: https://search.r-project.org/CRAN/refmans/esback/html/esr_backtest.html. The link also includes the details of the three versions of the backtests.. These ESR backtests (two-sided), labeled as A, S, and I respectively, are also employed to backtest the ES forecasts from the 11 competing models.
As in Table 8, the three backtests return quite consistent results, with the Semi-DCC models being not rejected. On the 2.5% probability level the models that get rejected by all three backtests are DCC-N, DCC-clik-N and DCC-t. On the 1% and 5% levels, the DCC-N and DCC-QML are rejected twice, respectively.
@@
$\bm{\alpha=1\%}$
CARE-SAV 22% 22% 21% 0
CARE-IG 100% 100% 87% 0
ES-CAViaR-SAV 89% 89% 86% 0
ES-CAViaR-IG 96% 96% 75% 0
DCC-QML 65% 64% 94% 0
DCC-N 2% 2% 7% 2
DCC-clik-QML 94% 94% 97% 0
DCC-clik-N 9% 9% 21% 0
DCC-t 10% 10% 7% 0
Semi-DCC-AL 99% 99% 87% 0
Semi-DCC-FZ0 99% 99% 86% 0
$\bm{\alpha=2.5\%}$
@@
Semi-DCC-AL 87% 87% 71% 0
Semi-DCC-FZ0 85% 85% 71% 0
$\bm{\alpha=5\%}$
CARE-SAV 18% 18% 12% 0
CARE-IG 16% 16% 10% 0
ES-CAViaR-SAV 17% 17% 12% 0
ES-CAViaR-IG 14% 14% 10% 0
DCC-QML 2% 2% 17% 2
DCC-N 27% 27% 41% 0
DCC-clik-QML 11% 12% 19% 0
DCC-clik-N 66% 65% 46% 0
DCC-t 38% 38% 52% 0
Semi-DCC-AL 43% 44% 21% 0
Semi-DCC-FZ0 47% 47% 25% 0
Lastly, comparing to the univariate semi-parametric risk forecasting models, such as ES-CAViaR or CARE, Semi-DCC models can be used for asset allocation. Therefore, in Appendix A we evaluate the out-of-sample performance of the Semi-DCC model in computing global minimum ES portfolios at the 2.5% level. Further, we compare the properties of the minimum ES portfolios obtained by Semi-DCC with those obtained by standard DCC models which are used as benchmarks. The effectiveness of the proposed Semi-DCC framework in portfolio optimization is demonstrated.
6 Conclusions
In this paper, we propose an innovative framework for forecasting portfolio tail risk in a multivariate and semi-parametric setting. The proposed framework is capable of modeling high dimensional return series parsimoniously and efficiently. Compared to the state-of-the-art univariate semi-parametric models, our approach delivers competitive risk forecasting performances and, in addition, it can be used for a wider range of applications including asset allocation. Compared to existing DCC estimation approaches, our empirical application shows that the Semi-DCC is able to deliver more accurate risk forecasts. Shifting the focus to the computation of minimum ES portfolios, we find that the Semi-DCC produces competitive performances compared to other DCC models, especially for the highly volatile “COVID 19” period.
Potential projects for future research include a deeper investigation of the performance of the Semi-DCC model in asset allocation, as well as an extension of the proposed framework to other settings such as the prediction of CoVaR (Adrian and Brunnermeier, 2016), building on recent results in semi-parametric estimation by Fissler and Hoga (2024) and Dimitriadis and Hoga (2024).
Gerlach and Wang (2020) extend the framework in Taylor (2019) by incorporating realized measures as exogenous variables, showing improved VaR and ES forecast accuracy. As our analysis in this paper is based on low-frequency daily data, our research plans also include extending the Semi-DCC model by incorporating high-frequency realized measures as exogenous variables to further improve its performance.
Furthermore, extending the Semi-DCC framework to more complex parameterizations, such as the AG-DCC model, is an interesting direction for future research. In this regard, it is worth noting that shifting the focus to alternative parameterizations could also pave the way for the investigation of alternative empirical frameworks. For example, a variant of the AG-DCC has recently been used by Pineda et al. (2022) to analyze financial contagion.
Finally, inspired by the results of our asset allocation exercise, we expect that Semi-DCC models could be well suited for tail risk hedging applications, especially when the hedged assets exhibit extreme volatility dynamics. However, despite the empirical relevance of this topic, its investigation goes beyond the scope of this paper and is currently left for future research.
Disclosure of Interest
No conflict of interests to be declared.
Disclosure of Funding
Data Availability Statement
Appendix A Minimum ES portfolios
We compare the properties of the 2.5% minimum ES portfolios obtained by Semi-DCC-AL with those obtained by standard DCC models, which are used here as benchmarks. Our interest for the computation of minimum ES portfolios is motivated by the prominent role that this risk measure is assuming in the international regulatory framework.
In order to compute minimum ES portfolios, we develop a two-stage procedure. At the initial stage, an estimate of the conditional variance and covariance matrix $H_{t}$ is obtained by fitting a Semi-DCC-AL model to an equally weighted portfolios. The optimal portfolio weights are then identified by minimizing the portfolio ES conditional on the first stage volatility estimates. There are two main reasons for our preference for this type of approach: first, it is computationally convenient as it avoids re-estimating the model at each iteration, and second, it can be easily implemented for semi-parametric models other than the Semi-DCC-AL model, thus facilitating comparison with benchmark DCC models.
@@ Appendix A Minimum ES portfolios
    Given an initial set of portfolio weights $\mathbf{w}$, fit a Semi-DCC-AL (or other standard DCC models) and save $\hat{H}_{t}$ for each in-sample time stamp $t$
@@
For all models out-of-sample 2.5% minimum ES portfolios are computed. Figure 5 illustrates the Semi-DCC-AL optimized minimum ES portfolio weights for all the forecasting steps. Table 9 compares the statistical properties of the minimum ES portfolio obtained by the Semi-DCC-AL model with the benchmarks. Firstly, we focus on the value of the empirical ES of optimized portfolios which is expected to be higher (closer to 0) for more effective risk minimization strategies. Compared to the EW portfolio, all estimated DCC models yield substantial reductions in terms of out-of-sample empirical ES. Meanwhile, Semi-DCC-AL and the other DCC models under analysis produce quite similar performances for the whole-out-of-sample period, with Semi-DCC-AL and DCC-clik-QML slightly outperforming the other DCC models. Such results lend evidence on the general validity of the proposed Semi-DCC-AL for asset allocation purposes. Regarding other descriptive statistics of portfolio returns, the Semi-DCC-AL produces smallest kurtosis and is least skewed. It also has the most conservative minimum value. These results provide evidence that the returns of the minimum ES portfolio obtained by the Semi-DCC-AL model are characterized by lighter tails and a less elongated left tail compared to the benchmarks, leading to a lower probability of occurrence of negative extreme events.
Figure 5: Boxplot of the Semi-DCC-AL 2.5% minimum ES portfolio weights for all forecasting steps.
Table 9: 2.5% minimum ES portfolio returns under different portfolio configurations produced by various models. Portfolio descriptive statistics include: $\widehat{ES}_{\alpha}$: sample $\alpha$-level ES ; $\kappa$: sample kurtosis; $skew$: sample skewness; $min$: sample minimum.
DCC-QML DCC-clik-QML DCC-t EW Semi-DCC-AL
@@
In addition, to examine the differences in the performance of the different models over a shorter time horizon, Figure 6 shows the moving empirical tail-mean (2.5% ES) of the optimised portfolio returns computed over a 250-day annual window. It is of particular interest to focus on the “COVID-19” period, which coincides with the negative peak in the plot, approximately centred on day 1000. First, we see that, in this highly risky period, the EW allocation is not effective in risk diversification since it yields remarkably lower ES values. At the same time, the DCC-t model, which generates competitive portfolio ES for the entire forecast period, is outperformed by all other semi-parametric DCC-type models for the high volatility “COVID-19” period, demonstrating the effectiveness of the semi-parametric approach, which is free of return distribution selection. Finally, focusing on the semi-parametric methods, it is evident that Semi-DCC-AL is able to produce preferred portfolio ES forecasts compared to DCC-QML and DCC-clik-QML for the “COVID-19” period.
@@
Figure 6: Moving 2.5% ES (sample tail mean) of optimized portfolio returns.